Determine the volume of the solid formed when the region enclosed by the curve y=√x, the x-axis, and the line x = 4, is revolved about the line x = 4.


1
Expert's answer
2021-05-24T02:28:02-0400

"V=\pi. \int _a^b f(x) ^2 dx"

"V= \pi. \int_0^4\sqrt x) ^2 dx"

"V=\pi. \int_0^4 x dx"

"V=\pi. [x^2/2]_0^4"

"V= \pi. [4^2/2] - \pi. [0^2/2]"

"V= 8\pi"

"V= 8 \u00d73.142"

"V=25.136" cubic units